0.5 2.25 14.0625 _______65534.0000152587
Answer by richwmiller(17219)

1/2, 9/4, 225/16
num 1^2 3^2 15^2 19^2 27^2
den 2 2^2 4^2 8^2 16^2
add numerator and denominator
1+2=3
3+2=5
15+4=19
but these don't work
19^2=361/64=45.5625
0.5 2.25 14.0625 1.41015625 65534.0000152587
27^2/16^2